# An Experimental Evaluation of Imputation Models for Spatial-Temporal Traffic Data

### v1

@@ IV-A Overview on Imputation Models for Time series
Li X et al. - ICDE’23 [33] BiSIM MTS imputation RNN + Attention unsupervised Indoor Positioning 0
@@
$\bigstar$ Li M et al. - ICDE’23 [71] PriSTI MTS imputation GNN + Attention + Diffusion Model self-supervised Air Quality / Traffic 4
Liu H et al. - VLDB’23 [50] HMTRL STG imputation GCN + GRU + Attention self-supervised Traffic 4
@@
Chen P et al. - TITS’24 [76] LRTC-SCAD MTS imputation TC unsupervised Traffic 0
Zeng Z et al. - TITS’24 [77] TNN-HTV MTS imputation TC unsupervised Traffic 2
Chen P et al. - TITS’24 [78] T-DAE STG imputation Transfer Learning GAN Traffic 2
Yang H et al. - TITS’24 [79] LFA-TRCE MTS imputation TC unsupervised Traffic 0
Shu H et al. - TITS’24 [80] LRTC-3DST STG imputation TC unsupervised Traffic 1
Wei X et al. - TITS’24 [81] SAGCIN STG imputation Attention + GCN self-supervised Traffic 0
Li B et al. - TITS’24 [82] CLRTR MTS imputation TC unsupervised Traffic 2
@@
Cheng S et al. - TITS’24 [72] FastSTI STG imputation Attention + GCN + Diffusion Model self-supervised Traffic 0
@@
Liu D et al. - TKDE’24 [83] SGMCAI-DiT MTS imputation Attention + Diffusion Model self-supervised Industry 5
Chen X et al. - TKDE’24 [84] LCR-2D MTS imputation TC unsupervised Traffic 17
@@
Park Byoungwoo et al. - NeurIPS’24 [85] PBDF MTS imputation Attention self-supervised Weather / Clinical 0
Obata Kohei et al. - KDD’24 [86] MissNet MTS imputation State-Space Model self-supervised Synthetic / Motion / Temperature 0
Nie T et al. - KDD’24 [55] ImputeFormer STG imputation Attention self-supervised Traffic 13
Zhang Z et al. - KDD’24 [87] LTVTI MTS imputation Attention self-supervised AIS 0
@@
IV-B Model Taxonomy for Traffic Data Imputation
To deepen our understanding of the similarities and differences among various imputation models for traffic data, we propose a practice-oriented taxonomy based on two criteria: spatial-temporal modeling techniques employed to capture the correlations in data, and loss design utilized for model training. Both criteria are closely tied to practical code implementation. Notably, an imputation model for traffic data may integrate a combination of these techniques and design its loss according to its specific model type and training strategy.
IV-B1 Spatial-Temporal Modeling Techniques
@@ V-B $E^{2}$
$\mathcal{L}_{D}=-D(\mathbf{X})+D(\hat{\mathbf{X}})$ (8)
@@
$\mathcal{L}_{G}=\lambda\|\mathbf{X}\odot\mathbf{M}-G(\mathbf{X}+\eta)\odot\mathbf{M}\|_{2}^{2}-D(\mathbf{X}^{\prime})$ (9)
@@ V-E LATC
where $d$ is a hyperparameter to determine the number of historical values, $h_{i}$ represents the time interval between the time of the $i$-th historical value and the current time $t$, =$\mathcal{H}=\{h_{1},\cdots,h_{d}\}$ denotes the set of these time intervals, and $\mathbf{W}\in\mathbb{R}^{N\times d}$ is the coefficient matrix that models the contribution of each historical value to the current time step.
Additionally, LATC preserves global trends in the imputed data by introducing a forward tensorization operator $\mathcal{Q}(\cdot)$ that transforms the imputed traffic matrix $\hat{\mathbf{X}}$ into a third-order tensor $\hat{\mathcal{X}}=\mathcal{Q}(\hat{\mathbf{X}})$. To ensure the low-rank characteristic of the tensor and capture the global patterns, LATC minimizes the nuclear norm of the third-order tensor $\|\hat{\mathcal{X}}\|_{r,*}$. Formally,
@@
[…]
where the operation $\mathcal{P}_{\Omega}(\cdot)$ retains the observable values while setting the missing positions to 00, and $r$ is a truncation parameter.
V-F PriSTI
@@ V-G GCASTN
$\mathcal{L}_{G}=\sum(\left\|\hat{\mathbf{X}}^{\prime}-{\mathbf{X}}\right\|_{2}^{2}+\left\|\hat{\mathbf{X}}^{\prime\prime}-{\mathbf{X}}\right\|_{2}^{2})$ (14)
@@
$\mathcal{L}_{C}=\sum(\left\|\hat{\mathbf{X}}^{\prime}-\hat{\mathbf{X}}^{\prime\prime}\right\|_{2}^{2}),$ (15)
where $\hat{\mathbf{X}}^{\prime}$ and $\hat{\mathbf{X}}^{\prime\prime}$ is the first and second positive sample. The total loss $\mathcal{L}=\mathcal{L}_{G}+\alpha\mathcal{L}_{C}$ by a weight $\alpha$.
V-H AGCRN
@@ V-J ImputeFormer
$\begin{split}\mathcal{L}_{\mathrm{recon}}&=\sum\|\mathbf{M}\odot(\widehat{\mathbf{X}}-\mathbf{X})\|_{1},\\ \mathcal{L}_{\mathrm{FIL}}&=\sum\|\mathrm{Flatten}(\mathrm{FFT}(\bar{\mathbf{X}})\|_{1},\\ \bar{\mathbf{X}}&=\mathbf{M}\odot\mathbf{X}+(\mathbf{1-M})\odot\widehat{\mathbf{X}},\end{split}$ (16)
@@
VI Experiments
VI-A Datasets
@@
VI-A1 Highway Traffic Flow
@@
VI-A2 Highway Traffic Speed
@@ VI-B Evaluation Pipeline
VI-B1 Missing Scenario Construction
@@ VI-B3 Model Construction
VI-B4 Performance Evaluation
We conduct a detailed evaluation of 20 missing scenarios and 10 baselines across three key aspects. For effectiveness, we use three commonly adopted metrics to assess mean error: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). For efficiency, we compare their training time, inference time, and memory usage. Additionally, we evaluate their imputation performance across both difficult and simple time intervals to simulate varying real-world traffic conditions. Note that our evaluation only focuses on the performance at the locations where data is missing.
VI-C Performance Evaluation Results
VI-C1 Imputation Performance on Different Datasets
@@
Figure 4: Performance of traffic data imputation for different models with 20 missing scenarios at PEMS04.
Figure 5: Performance of traffic data imputation for different models with 20 missing scenarios at PEMS08.
As shown in Figures 5, 5, and 6, we observe that the performance declines as the missing rate increases. When the missing rate is below 0.5, the performance change of each model remains relatively stable. However, as the missing rate reaches 0.7 and 0.9, the performance of all models significantly deteriorates. This indicates that the noise of missing values at lower missing rates is minimal and does not substantially impact the model. In contrast, when the missing rate exceeds 0.5, the abundance of missing values disrupts the inherent spatial-temporal relationships in the data.
Figure 6: Performance of traffic data imputation for different models with 20 missing scenarios at Seattle.
ImputeFormer, GCASTN, BRITS, and PriSTI are the top-4 best deep-learning models for traffic data imputation, a common characteristic of their specialized design for handling missing data. ImputeFormer incorporates low-rankness into Transformers to capture spatiotemporal structures and mitigate the noise impact of missing value. GCASTN and BRITS use the time decay strategy to introduce time intervals between observed and missing values, while PriSTI applies linear imputation to capture temporal context. Therefore, efficiently incorporating prior knowledge of missing data may be a promising direction in traffic data imputation.
@@
TABLE III: Performance comparison at difficult and simple intervals.
Model Difficult Intervals Simple Intervals
@@
VI-C2 Performance During Difficult and Simple Intervals
Most imputation models evaluate average performance in the test set, which fails to capture their strengths and weaknesses across different intervals. For instance, data imputation is more challenging during periods of fluctuating flow changes, while smoother flow transitions are easier to handle. To address this, we sort test samples by variance, with the top 25% with the highest variance classified as difficult intervals and the bottom 25% as simple intervals. We report performance across these intervals using the PEMS04 dataset with a 0.5 missing rate under the SRTR missing pattern, as shown in Table III. The results show that performance decreases significantly in difficult intervals due to frequent fluctuations, whereas simple intervals, characterized by smoother flow, exhibit more predictable temporal relationships.
By comparing models, ImputeFormer and GCASTN demonstrate superior performance in difficult and simple intervals, respectively, highlighting the importance of spatial-temporal modeling. ImputeFormer effectively captures global knowledge and adapts to fluctuations, while GCASTN is well-suited for steady trends by utilizing the time decay strategy. GCASTN also achieves the best and second-best performance in difficult intervals due to its contrastive learning strategy, which enhances robustness. BRITS and E2GAN underperform as they focus solely on temporal dependencies, while spatial-temporal modeling methods such as ASTGNN and AGCRN produce more balanced performance.
Figure 10: Visualization of model performance across different missing patterns with the missing rate of 0.5 on PEMS04.
VI-C3 Ranks on Model Performance across Different Missing Patterns
The different missing patterns caused the imputation performances of the model to be different. To explore which technology is effective for missing patterns, we count the times of each model that achieved the best, the top-2, and the top-3 performance. If a model has more times, its performance is better. As shown in Figure 7, We observe that only BRITS, GCASTN, PriSTI, ImputeFormer, and LATC achieve the top 3 best performance. A common characteristic of them is introducing prior knowledge, where BRITS, GCASTN, and PriSTI use time delay, and ImputeFormer and LATC consider low-rank characteristics to capture the global correlation. Yet other models only consider the spatio-temporal correlation, failing to efficiently imputation since extra noise of missing data leads to the spatiotemporal correlation being inaccurate. Thus, we guess efficiently extra prior knowledge from data may be a promising direction for more accurate imputation.
TABLE IV: Time and memory cost. / notes that the model does not need to infer stage.

### v2

@@ III-A4 Time Series Probabilistic Imputation
Li et al. - ICDE’23 [44] BiSIM MTS imputation RNN + Attention Predictive (RMSE) Reconstruction-based SSL Indoor Positioning
@@
$\bigstar$ Li et al. - ICDE’23 [84] PriSTI MTS imputation GNN + Attention Generative (Diffusion) Masked SSL Air Quality / Traffic
Liu et al. - VLDB’23 [61] HMTRL STG imputation GNN + RNN + Attention Predictive Masked SSL Traffic
Li et al. - SIGMOD’23 [89] SSIN ST Kriging Attention Predictive Masked SSL Raingauge
@@
Chen et al. - TITS’24 [90] LRTC-SCAD MTS imputation TC Predictive Reconstruction SSL Traffic
Zeng et al. - TITS’24 [91] TNN-HTV MTS imputation TC Predictive Reconstruction SSL Traffic
Yang et al. - TITS’24 [92] LFA-TRCE MTS imputation TC predictive Reconstruction SSL Traffic
Shu et al. - TITS’24 [93] LRTC-3DST STG imputation TC predictive Reconstruction SSL Traffic
Wei et al. - TITS’24 [94] SAGCIN STG imputation Attention + GCN Predictive(MAE) Masked SSL Traffic
Li et al. - TITS’24 [95] CLRTR MTS imputation TC predictive Reconstruction SSL Traffic
@@
Cheng S et al. - TITS’24 [85] FastSTI STG imputation Attention + GCN + Diffusion Model Generative (Diffusion) Masked SSL Traffic
@@
Liu et al. - TKDE’24 [96] SGMCAI-DiT MTS imputation Attention + Diffusion Model Genertaive (Diffusion) Masked SSL Industry
Chen et al. - TKDE’24 [97] LCR-2D MTS imputation TC Predictive Reconstruction SSL Traffic
@@
Park Byoungwoo et al. - NeurIPS’24 [98] PBDF MTS imputation Attention Predictive(MSE) Masked SSL Weather / Clinical
Deng et al. - NeurIPS’24 [99] OPCR STG imputation GCN + Attention Predictive(MAE) Masked SSL Traffic
$\bigstar$ Nie et al. - KDD’24 [66] ImputeFormer STG imputation Attention Predictive(MAE) Masked SSL Traffic
Zhang et al. - KDD’24 [100] LTVTI MTS imputation Attention+Diffusion Generative (Diffusion) Masked SSL AIS
Wang et al. - IJCAI’25 [101] STAMImputer STG imputation GCN + Attention + MoE Predictive(MAE) Masked SSL Traffic
Zheng et al. - TITS’25 [102] PRMAN STG imputation Attention Predictive(MSE) Masked SSL Traffic
Yang et al. - AAAI’25 [103] GSLI STG imputation GCN + Attention Predictive(MSE) Masked SSL Traffic
$\bigstar$ Huang et al. - AAAI’25 [104] STD-PLM STG imputation GCN + Attention Predictive(MAE) Masked SSL Traffic
Xu et al. - AAAI’25 [105] KITS ST Kriging GCN Predictive(MAE) Masked SSL Traffic / Air Quality
Liu et al. - TKDE’25 [106] TIDER MTS imputation TC Predictive (MSE) Reconstruction SSL Traffic / Solar power
III-B Model Taxonomy for Traffic Data Imputation
To deepen our understanding of the relationships and differences among imputation models applicable to traffic data, we propose a practice-oriented taxonomy based on two key dimensions: spatial-temporal modeling techniques, which capture inherent correlations in the data, and loss design, which guides model training and is tailored to the specific model type and training strategy, as shown in Figure 4. Both dimensions are closely aligned with practical implementation considerations. Importantly, a traffic data imputation model may combine multiple modeling techniques and adapt its loss function to suit its specific design and training paradigm. Building on this taxonomy, together with the four categories of time-series imputation tasks discussed above, we provide a comprehensive summary of representative works in Table I.
@@ III-B Model Taxonomy for Traffic Data Imputation
Figure 4: The practice-oriented taxonomy on imputation models applicable to traffic data.
III-B1 Spatial-Temporal Modeling Techniques
@@ IV-B $E^{2}$GAN
$D(𝐗^)\mathcal{L}_{D}$=-D($\mathbf${X})+D($\hat$$\mathbf${X}}) (8)
@@
[…]
@@ IV-E LATC
where $d$ is a hyperparameter to determine the number of historical values, $h_{i}$ represents the time interval between the time of the $i$-th historical value and the current time $t$, $\mathcal{H}=\{h_{1},\cdots,h_{d}\}$ denotes the set of these time intervals, and $\mathbf{W}\in\mathbb{R}^{N\times d}$ is the coefficient matrix that models the contribution of each historical value to the current time step.
[…]
@@
$\text{s.t.}\text{ }\hat{\mathcal{X}}=\mathcal{Q}(\hat{\mathbf{X}}),\mathcal{P}_{\Omega}(\hat{\mathbf{X}})=\mathcal{P}_{\Omega}(\mathbf{X}),$
where the operation $\mathcal{P}_{\Omega}(\cdot)$ retains the observable values while setting the missing positions to 0, and $r$ is a truncation parameter.
IV-F PriSTI
@@ IV-G GCASTN
$\mathcal{L}_{G}=\sum(\left\|\hat{\mathbf{X}}_{1}-{\mathbf{X}}\right\|_{2}^{2}+\left\|\hat{\mathbf{X}}_{2}-{\mathbf{X}}\right\|_{2}^{2})$ (14)
@@
$\mathcal{L}_{C}=\sum(\left\|\hat{\mathbf{X}}_{1}-\hat{\mathbf{X}}_{2}\right\|_{2}^{2}),$ (15)
where $\hat{\mathbf{X}}_{1}$ and $\hat{\mathbf{X}}_{2}$ are two positive data views. The total loss $\mathcal{L}=\mathcal{L}_{G}+\alpha\mathcal{L}_{C}$ by a weight $\alpha$.
IV-H AGCRN
@@ IV-J ImputeFormer
$\begin{split}\mathcal{L}_{\mathrm{recon}}&=\sum\|\mathbf{M}\odot(\widehat{\mathbf{X}}-\mathbf{X})\|_{1},\\ \bar{\mathbf{X}}&=\mathbf{M}\odot\mathbf{X}+(\mathbf{1-M})\odot\widehat{\mathbf{X}},\\ \mathcal{L}_{\mathrm{FIL}}&=\sum\|\mathrm{Flatten}(\mathrm{FFT}(\bar{\mathbf{X}})\|_{1},\end{split}$ (16)
@@
IV-K STD-PLM
STD-PLM [104] is a unified spatial-temporal prediction and imputation model built upon pre-trained language models. By converting spatial-temporal data into token sequences through a specially designed spatial-temporal tokenizer, STD-PLM leverages the powerful comprehension and reasoning capabilities of pre-trained language models for both forecasting and imputation tasks. To enhance the model’s awareness of the spatial topological structure inherent in spatial-temporal data, STD-PLM introduces an adaptive node embedding mechanism based on the adjacency matrix and a region-level token construction strategy. These designs jointly enable STD-PLM to achieve competitive performance on both prediction and imputation benchmarks.
Loss design: Following PriSTI, STD-PLM adopts a masked SSL strategy, wherein additional conditional missingness is artificially injected into the training data to construct supervision samples.
V Experiments
V-A Datasets
@@
V-A1 Highway Traffic Flow
@@
V-A2 Highway Traffic Speed
@@
V-A3 TW
We also collected a new dataset TW from public sources for our experiments. The TW dataset comprises traffic flow data from 315 sensors across Taiwan, recorded at 5-minute intervals, spanning the period from November 1st, 2020 to November 21st, 2020.
@@
Dataset Spatiotemporal Correlation Analysis. As shown in Figure 12(a), we conduct a dataset-level macro spatiotemporal correlation analysis. For temporal correlation, we consider the traffic flow/speed for each station as a time series and calculate the dynamic time warping (DTW) similarity between time series from one day to the following day. For spatial correlation, we use traffic flow data for one day at each node, compute the DTW across $N$ sequences, and analyze their similarity. The experimental results are illustrated in the figure. We observe that PEMS04 and PEMS08 show stronger temporal correlations, with 95.11% and 96.47% of the data reaching a temporal similarity of 80%. In contrast, only 7.30% of TW’s data reached 80% similarity, and the Seattle dataset only has 4.64% data.
As for spatial correlation, PEMS04 and PEMS08 demonstrate stronger results, reaching 80% similar with 62.31% and 72.62% data, respectively. In Seattle, only 1.30% of nodes achieve 80% similarity. In TW, 8.29% of nodes reached 80% similarity.
(a) Spatiotemporal Correlation of different datasets.
@@ V-B Evaluation Pipeline
V-B1 Missing Scenario Construction
@@ V-B3 Model Construction
V-B4 Performance Evaluation
We conduct a detailed evaluation of 20 missing scenarios and 10 baselines across three key aspects. For effectiveness, we use three commonly adopted metrics to assess mean error: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). For efficiency, we compare their training time, inference time, and memory usage. Additionally, we evaluate their imputation performance across both challenging and stable periods to simulate varying real-world traffic conditions. Note that our evaluation only focuses on the performance at the locations where data is missing.
V-C Effectiveness and Robustness Evaluations
V-C1 Imputation Performance on Different Datasets
@@
Figure 13: Performance of traffic data imputation for different models under 20 missing scenarios at PEMS04.
Figure 14: Performance of traffic data imputation for different models under 20 missing scenarios at PEMS08.
As shown in Figures 14, 14, 15, and 16, we observe that the performance declines as the missing rate increases. When the missing rate is below 0.5, the performance change of each model remains relatively stable. However, as the missing rate reaches 0.7 and 0.9, the performance of all models significantly deteriorates. This indicates that the noise of missing values at lower missing rates is minimal and does not substantially impact the model. In contrast, when the missing rate exceeds 0.5, the abundance of missing values disrupts the inherent spatial-temporal relationships in the data.
Figure 15: Performance of traffic data imputation for different models under 20 missing scenarios at Seattle.
Figure 16: Performance of traffic data imputation for different models under 20 missing scenarios at TW.
ImputeFormer, GCASTN, BRITS, and STD-PLM are the top-4 best deep-learning models for traffic data imputation, a common characteristic of their specialized design for handling missing data. ImputeFormer incorporates low-rankness into Transformers to capture spatiotemporal structures and mitigate the noise impact of missing value. GCASTN and BRITS use the time decay strategy to introduce time intervals between observed and missing values, while STD-PLM directly learns the spatial relationship of sensors by introducing the prior of spatial topology. Therefore, efficiently incorporating prior knowledge of missing data may be a promising direction in traffic data imputation.
@@
Figure 17: Comparison of models under different scenarios with low missing rates.
@@
Figure 18: Comparison of models under different scenarios with high missing rates.
V-C2 Performance under Different Missing Scenarios
To comprehensively evaluate the overall performance of various models under different missing data scenarios, we categorize the missing rate into two levels: low missing rate ($50\%$) and high missing rate ($50\%$). Based on this categorization, we plot corresponding radar charts for each dataset and each evaluation metric to visualize model performance across different missing data settings, as shown in Figure 17 and 18.
To make performance comparisons more intuitive, we select some of the well-performing models and normalize their metrics to a scale of [0, 100], where higher values indicate better results. Under this setting, the area enclosed by a model reflects its overall effectiveness, with larger areas representing stronger performance. Notably, the connections between two missing patterns on the radar charts also provide insights into a model’s capacity to handle hybrid missing patterns. For example, the line connecting SRTR and SCTR suggests that the missing mode is the mixing missing of both SRTR and SCTR.
We observe that dataset characteristics and missing data patterns significantly affect model performance. On the PEMS04 and PEMS08 datasets, ImputeFormer performs best under SR- missing patterns, while LATC shows superior results in SC- missing patterns. Both models share the use of tensor decomposition strategies, indicating that tensor decomposition is highly effective for handling missing data. Additionally, GCASTN and STD-PLM also demonstrate comparable performance.
For the Seattle traffic speed dataset, BRITS consistently outperformed other methods, and GCASTN is the second-best-performing model. This may be attributed to the relatively stable nature of speed over time, where temporal fluctuations are minimal. In such cases, a simple RNN with a time decay mechanism can effectively model the data, whereas more complex models might lead to performance degradation.
On the TW dataset, IGNNK achieves the best performance in most missing patterns, especially under TC- missing. This is mainly because the adjacency matrix of the TW dataset is relatively sparse, and the sensor values are similar to those of their neighboring sensors. Therefore, a simple GNN is sufficient to model the data effectively. However, under the SCTR missing pattern, IGNNK performed poorly due to the unavailability of signals from many nearby spatial sensors. In contrast, LATC considers global spatial relationships and achieves the best results in this case. Under the SCTC missing pattern, where most information is unavailable, models that rely on complex spatiotemporal relationships tend to perform worse.
We also count how many times each model achieves the best, top-2, and top-3 performance. A higher count indicates stronger overall robustness and adaptability. As shown in Figure 19, We observe that BRITS, GCASTN, PriSTI, ImputeFormer, and LATC achieve the top 3 best performances. A common characteristic of these models is the introduction of prior knowledge, where BRITS, GCASTN, and PriSTI utilize time delay, and ImputeFormer and LATC consider low-rank characteristics to capture global correlation. Yet other models only consider the spatio-temporal correlation, failing to efficiently impute, since extra noise of missing data leads to the spatiotemporal correlation being inaccurate. Thus, we guess that efficiently extracting extra prior knowledge from data may be a promising direction for more accurate imputation.
(a) Visualization of model performance across different missing patterns with the missing rate of 0.5 on PEMS04.
TABLE III: Performance comparison during challenging periods and stable periods.
Model Challenging Periods Stable Periods
@@ V-C2 Performance under Different Missing Scenarios
STD-PLM 23.57 35.38 9.14 7.64 12.00 18.63
V-C3 Performance During Challenging and Stable Periods
Most imputation models evaluate average performance in the test set, which fails to capture their strengths and weaknesses across different intervals. For instance, data imputation is more challenging during periods of fluctuating flow changes, while smoother flow transitions are easier to handle. To address this, we sort test samples by variance, with the top 25% with the highest variance classified as challenging periods and the bottom 25% as stable periods. We report performance across these periods using the PEMS04 dataset with a 0.5 missing rate under the SRTR missing pattern, as shown in Table III. The results show that performance decreases significantly during challenging periods due to frequent fluctuations, whereas stable periods, characterized by smoother flow, exhibit more predictable temporal relationships.
By comparing models, GCASTN, STD-PLM, and ImputeFormer demonstrate superior performance in challenging and stable periods, respectively, highlighting the importance of spatial-temporal modeling. ImputeFormer effectively captures global knowledge and adapts to fluctuations. STD-PLM separately models the spatiotemporal relationships between nodes and leverages the language reasoning capabilities of LLM to learn traffic patterns in both challenging and stable periods, while GCASTN is well-suited for steady trends by utilizing the time decay strategy. BRITS and E2GAN underperform as they focus solely on temporal dependencies, while spatial-temporal modeling methods such as ASTGNN and AGCRN produce more balanced performance.
TABLE IV: Time and memory cost. / notes that the model does not need an inference stage.